OpenStudy (kendricklamar2014):

What is an equation for the line with slope 2/3 and y-intercept 9? A. Y = 2/3x B. 9x C. Y = 2/3x + 9 D. 9x + 2/3

OpenStudy (kendricklamar2014):

@TheSmartOne @StudyGurl14

TheSmartOne (thesmartone):

y=mx+b where m is slope and b is y-intercept.

TheSmartOne (thesmartone):

So just plug in m=2/3 and b=9

OpenStudy (kendricklamar2014):

So y= 2/3 + 9

TheSmartOne (thesmartone):

2/3x**

TheSmartOne (thesmartone):

So it is C.

OpenStudy (kendricklamar2014):

Thanks

OpenStudy (kendricklamar2014):

100% thanks @TheSmartOne

TheSmartOne (thesmartone):

No prob. :) I don't suppose you went on another potty break @StudyGurl14

OpenStudy (studygurl14):

Nope. Openstudy this time @TheSmartOne